From past experience, a company has found that in a box of four triple-A batteries, 74% contain no defective battery, 7% contain one defective battery, 4% contain two defective batteries, 6% contain three defective batteries, and 9% contain all defective batteries. Calculate the mean, variance, and standard deviation for the defective batteries


1
Expert's answer
2022-04-08T04:43:22-0400

Let X - the random variable of the number of defective batteries in a box.


The mean:

"\mu=\sum x_i\cdot P(x_i)=\\
=0\cdot0.74+1\cdot0.07+2\cdot0.04+\\
+3\cdot0.06+4\cdot0.09=0.69."


The variance:

"\sigma^2=\sum(x_i-\mu)^2\cdot P(x_i),"

"X-\mu=\\
=\{ 0-0.69, 1-0.69, 2-0.69, 3-0.69,4-0.69\}=\\
=\{-0.69,0.31,1.31,2.31,3.31\},"

"\sigma^2=(-0.69)^2\cdot0.74+0.31^2\cdot0.07+1.31^2\cdot0.04+\\
+2.31^2\cdot0.06+3.31^2\cdot0.09=1.734."


The standard deviation:

"\sigma=\sqrt{1.734}=1.317."